# Shifting Attention to You: Personalized Brain-Inspired AI Models

### v1

@@ Abstract
The integration of human and artificial intelligence represents a scientific opportunity to advance our understanding of information processing, as each system offers unique computational insights that can enhance and inform the other. The synthesis of human cognitive principles with artificial intelligence has the potential to produce more interpretable and functionally aligned computational models, while simultaneously providing a formal framework for investigating the neural mechanisms underlying perception, learning, and decision-making through systematic model comparisons and representational analyses. In this study, we introduce personalized brain-inspired modeling that integrates human behavioral embeddings and neural data to align with cognitive processes. We took a stepwise approach, fine-tuning the Contrastive Language–Image Pre-training (CLIP) model with large-scale behavioral decisions, group-level neural data, and finally, participant-level neural data within a broader framework that we have named CLIP-Human-Based Analysis (CLIP-HBA). We found that fine-tuning on behavioral data enhances its ability to predict human similarity judgments while indirectly aligning it with dynamic representations captured via magnetoencephalography (MEG). To further gain mechanistic insights into the temporal evolution of cognitive processes, we introduced a model specifically fine-tuned on millisecond-level MEG neural dynamics (CLIP-HBA-MEG). This model resulted in enhanced temporal alignment with human neural processing while still showing improvement on behavioral alignment. Finally, we trained individualized models on participant-specific neural data, effectively capturing individualized neural dynamics and highlighting the potential for personalized AI systems. These personalized systems have far-reaching implications for the fields of medicine, cognitive research, human-computer interfaces, and AI development.
@@ 1 Introduction
Artificial intelligence (AI) systems are fundamentally composed of four key components: the model architecture that structures the processing mechanism, the learning rules that guide parameter updates, the objective functions that shape the learning path, and the datasets on which they are trained (Goodfellow et al., 2016; Saxe et al., 2021). While considerable progress has been made in each of these components and created much improved AI systems, aligning AI with human cognition remains a significant challenge (Ferrante et al., 2024; Lu et al., 2024; Rane et al., 2024). Such alignment holds the promise of not only enhancing the interpretability and robustness of AI models but also of deepening our understanding of human cognitive processes (Mahner et al., 2024).
Humans often rely on semantic abstraction and operate across varying levels of semantic categorization, making it essential for AI models to excel in both superordinate and fine-grained tasks (Goldstone, 1994; Schulte Im Walde and Frassinelli, 2022; Born et al., 2024). However, the representational strategies employed by humans and those utilized by AIs differ, highlighting the need for more sophisticated approaches to bridge this gap (Mahner et al., 2024). For instance, images that are indistinguishable to models–which form model metamers in later stages of neural networks–often appear obviously different to humans, demonstrating the presence of idiosyncratic, model-specific invariances that differ from the organization and representations of human sensory systems (Feather et al., 2023).
@@
While the shared architectural principles between deep neural networks (DNNs) and human cortical processing (Cichy et al., 2016) play a role in brain alignment, recent studies suggest that the path to better human alignment may lie more in how models are trained than in their fundamental architecture (Muttenthaler et al., 2023a). The choice of training components, such as the objective function, significantly impacts the learning of human-like invariances (Nanda et al., 2023). Thus, one way to make models more human is to incorporate human inductive biases into AI systems to mirror cognitive processes. Indeed, leveraging these biases through Bayesian model selection and contrastive learning has been shown to enhance human alignment and machine learning performance (Sinz et al., 2019; Marjieh et al., 2023, 2024).
In this study, we aim to enhance the alignment between AIs and human cognition by fine-tuning the CLIP model (Radford et al., 2021) using human behavioral embeddings and neural data. Notably, the behavioral embeddings, derived from large-scale behavioral modeling (Hebart et al., 2019, 2023), provide interpretable dimensions that reflect human mental representations of objects. By training the model to incorporate this human-derived mental embedding space, we enable it to predict human behavior and neural representations more accurately, while also achieving significantly enhanced performance on alignment benchmarks (Fu et al., 2023).
While large-scale behavioral embeddings offer a fixed output of generalizable perceptual representations, neural data provides more dynamic insights that can reflect the millisecond-level temporal evolution of human brain representations (Ruffle et al., 2024). Moreover, neural data has the potential to better capture individual variability across diverse populations where behavior may be challenging to collect (e.g., infants, non-verbal population) (Kucyi et al., 2024a). To explore this, we further introduced a dynamic neural fine-tuning process using MEG signals (CLIP-HBA-MEG) to better align the model’s temporal representations with the rapidly changing activity patterns of the human brain.
Through training with group-level MEG representations (Hebart et al., 2023), we observed that the model generalizes well across external datasets, encompassing various participants, conceptual stimuli, and visual conditions. This finding reinforced our confidence in the robustness and generalizability of the approach, allowing us to move forward to individualized perceptual fine-tuning with participant-specific neural data and capture unique neural dynamics and perceptual differences. Our approach holds promise for creating personalized AI systems that reflect distinct and individual cognitive patterns, bridging the gap between artificial and biological cognition. By embedding these individualized patterns into AI models, we enhance their interpretability and alignment while deepening our understanding of the interplay between minds and machines.
@@ 2 Results
Current AI models trained on large-scale image datasets develop inductive biases based on the statistical patterns within those datasets. However, a significant gap exists between machine-derived inductive biases and the way humans develop theirs through perceptual experience (Goyal and Bengio, 2022). To address this, we introduced human-like inductive biases to the CLIP Model in the form of 66 Sparse Positive Similarity Embedding (SPoSE) dimensions, derived from large-scale behavioral studies (Zheng et al., 2019). These dimensions have been shown to effectively predict human similarity judgments for object distinctions (Hebart et al., 2020).
We developed CLIP-HBA-Behavior (Figure 1) by fine-tuning the CLIP model, using the 66 SPoSE dimensions as textual inputs while simultaneously feeding THINGS images (Hebart et al., 2019) as visual inputs. The CLIP architecture encodes both modalities—textual and visual—separately, before binding them via dot product similarity. This mechanism evaluates the encoded visual features against each of the 66 dimensions, resulting in a 66-dimensional embedding representation for each image input.
@@ 2.1 Fine-Tuning CLIP with Human Behavioral Embeddings Enhances Alignment
As shown in Figure 2A, we evaluated the behaviorally fine-tuned model’s performance on a set of 48 object images from the THINGS database, which were fully sampled in a behavioral odd-one-out task (Hebart et al., 2020). These 48 objects were excluded from training to prevent data leakage. To assess the alignment between the model’s predictions and human behavior, we calculated the Spearman rank correlation ($\rho$) between the model’s predicted Representational Dissimilarity Matrix (RDM) (Kriegeskorte, 2008) and the ground-truth behavioral RDM. The behaviorally fine-tuned CLIP-HBA-Behavior achieved a Spearman correlation of 0.78 ($p<10^{-229}$) and a 95% confidence interval (CI) of [0.75, 0.80].
@@ 2.2 Improved Performance on Alignment Benchmark with CLIP-HBA-Behavior
To test our model’s robustness and performance, we benchmarked it against a state-of-the-art human visual similarity judgment metric—NIGHTS (Novel Image Generations with Human-Tested Similarity), a benchmark of 20,019 image triplets with human-tested perceptual similarity scores (Fu et al., 2023).
@@ 2.3 Improved Neural Alignment with CLIP-HBA-Behavior
While the enhanced behavioral correlation of our model was anticipated due to its training on behavioral data, understanding whether the model’s embedding space truly becomes more human-like requires evaluating how training exclusively on behavioral data impacts its alignment with neural representations. Therefore, we further evaluated the alignment between the model’s representations against neural data obtained from MEG recordings (Hebart et al., 2023).
@@
  • Colored Images without Background (Figure 2B, top right): Clear object stimuli with backgrounds featuring humans, animals, fruits, and man-made objects (Grootswagers et al., 2017a).
  • Monochromatic Clear Images (Figure 2B, bottom left): Clear, monochromatic images of animate and inanimate objects (Grootswagers et al., 2017b).
  • Monochromatic Images with Blur (Figure 2B, bottom right): Blurry, monochromatic images of animate and inanimate objects (Grootswagers et al., 2017b).
@@ 2.4 Dynamic Neural Fine-Tuning Captures Millisecond-Level Neural Dynamics
Building on the success of perceptual fine-tuning with behavioral data, which are large-scale, generalizable, and collected across many trials and participants, we sought to explore whether the model can improve by learning directly from human neural representations. Although typically collected from smaller participant samples, neural data can often be better suited to capture diverse neural profiles across different populations (Ye and Pandarinath, 2021). While behavioral embeddings provide static and generalizable perceptual representations, neural data offer dynamic insights, reflecting millisecond-level temporal patterns of human brain activity that cannot be captured through behavioral measures alone. Furthermore, neural data have the potential to better account for individual variability, providing a more personalized lens into how visual stimuli are processed (Pagan et al., 2024). To explore these possibilities, we developed a dynamic neural fine-tuning process using MEG signals, as demonstrated in Figure 3. The dynamically fine-tuned model, referred to as CLIP-HBA-MEG, adapts its visual stimulus embeddings to align with distinct stages of visual processing and the emergence of semantic representations.
@@ 2.6 Viewable Dynamic Attention of CLIP-HBA-MEG
Leveraging the dynamic embedding spaces of the neurally aligned CLIP-HBA-MEG, we can visualize the specific regions and pixels the model prioritizes to produce human-like neural responses at a millisecond resolution. By combining randomized input sampling of the visual stimuli (Petsiuk et al., 2018) (Kaniuth et al., 2024a) with the model’s dynamically learned visual scaler, we can quantify both where and to what extent the model attends within any image following stimulus onset. As shown in Figure 5, the highlighted regions indicate where the model predicts human observers would direct their attention at a given time point.
@@ 2.7 Individualized Models Capture Participant-Specific Neural Dynamics
As shown in Figure 6, we extended the dynamic neural fine-tuning pipeline of CLIP-HBA-Dynamic to train 15 personalized models, each fine-tuned on the individual MEG data of a single participant (Cichy et al., 2016). These personalized models achieved enhanced neural alignment specific to their respective participants, exhibiting sustained higher temporal alignment compared to the lower-bound noise ceiling, starting around 200 ms after stimulus onset.
@@ 3.1 Scalable and Transferable Applications
Building on these encouraging findings, our first area of focus was the scalability and transferability of the CLIP-HBA framework. CLIP-HBA-MEG leverages dynamic neural signals to capture real-time perceptual processes and higher-order conceptual organization, offering a nuanced window into human perception that static models cannot provide. In contrast to models that focus primarily on behavioral proxies—such as those that linearly map behavioral embeddings into model output (Kaniuth et al., 2024b) or perform generalizable local and global transforms to predict human similarity ratings (Muttenthaler et al., 2023b)—CLIP-HBA-MEG establishes a direct relationship with neural activity by explicitly capturing the emergence of information from dynamic neural signals. Recent research has shown that even when behavioral outcomes converge, the underlying neural dynamics can differ substantially between individuals (Renart and Machens, 2014; Bansal et al., 2021). Such findings underscore the critical importance of capturing dynamic neural activity. While earlier models—such as ReAlnet (Lu et al., 2024)—attempt to capture cross-modal EEG-fMRI alignments, they lack the temporal resolution necessary to track moment-to-moment perceptual shifts. Moreover, the CLIP-HBA framework can be fine-tuned sequentially, influencing the entire stream of information processing, which uniquely positions it as both a robust model of human perception and a versatile platform that can integrate diverse types of human behavior (e.g., MEG, MRI, behavioral data, electrophysiology). This adaptability is particularly useful in neuroscience applications, where neural data is often limited in scale.
@@ 3.2 Explainable Human-Driven AI Systems
In addition to capturing real-time perceptual processes, CLIP-HBA provides fully interpretable dimensions within the model embedding space (Hebart et al., 2020) that are absent in other models designed to incorporate human inductive biases (Lu et al., 2024). By integrating this embedding space with dynamic saliency mapping, we offer millisecond-level insights into both the model’s visual focus and the organization of its internal representations. Furthermore, findings indicate that when model decisions are transparent and systematically quantifiable, it becomes easier to evaluate the consistency of predictions, identify potential biases, and isolate errors for correction (Doshi-Velez and Kim, 2017; Rudin, 2019). These results support the view that interpretability is essential in applications where accurate, accountable decision-making is critical. Together, these advancements move AI closer to human-like cognition by promoting systems whose internal operations are more accessible and amenable to rigorous oversight (Zador et al., 2023; Sinz et al., 2019).
@@ 3.3 Expandable Architectures and Modalities for Future Directions
While our study demonstrates the effectiveness of fine-tuning CLIP to align AI models with human representations using behavioral and MEG data, our approach is not inherently limited to a specific model architecture, type of neural data, or sensory modality. The key factors contributing to the success of our method are the carefully designed loss functions and the modified mechanisms, such as feature reweighting (Kaniuth and Hebart, 2022) and noise injections. These techniques are broadly applicable to other transformer-based and deep neural network architectures, ensuring that our framework remains adaptable beyond CLIP. This allows our approach to be well-suited for expanding to additional sensory modalities. While our current work focuses on vision, future research could integrate other modalities such as auditory representations for multisensory learning (Baltrušaitis et al., 2018). By fine-tuning models with neural and behavioral data across multiple sensory domains, we can take the first steps toward a truly multisensory AI system—one that more closely mimics the way the human brain integrates information from diverse sensory inputs (Murray and Wallace, 2011; Tovar et al., 2020a). This expansion would not only enhance AI’s ability to model human perception but also open new possibilities for multisensory cognitive research and human-AI interaction.
@@ 3.4 Applications of Personalized Models
With the framework’s adaptability established, we now turn to its potential for personalization—a critical aspect of aligning AI with individual human cognition. The CLIP-HBA framework represents a significant advancement in personalized AI by integrating human behavioral embeddings with neural data to tailor models to individual cognitive profiles. This approach is deeply rooted in recent cognitive neuroscience research, which suggests that aligning artificial neural networks with biological cognition can enhance both interpretability and robustness (Doerig et al., 2023). By fine-tuning on millisecond-level MEG signals, the CLIP-HBA-MEG variant directly optimizes neural alignment, reflecting the neuroconnectionist framework’s emphasis on incorporating biological constraints rather than solely focusing on machine-centric performance. Moreover, the framework’s dynamic adaptation of learned representations based on neural feedback mirrors the functionality of digital twin systems (Jiang et al., 2021)—widely used in industrial and medical applications to model dynamic systems with real-time sensor data—thus acting as a cognitive digital twin that adjusts to an individual’s unique perceptual and neural idiosyncrasies.
Beyond the realm of neuroscience, personalized AI models like CLIP-HBA are making transformative strides in various domains, including education, healthcare, and industry. Advances in AI-driven learning style detection have demonstrated that machine learning can accurately classify users into distinct cognitive subtypes—such as visual, auditory, or kinesthetic learners (Kanchon et al., 2024)—which aligns with CLIP-HBA’s ability to fine-tune representations based on individual neural responses. Such personalization is critical for applications where individual differences matter; for example, in clinical contexts, these models can be harnessed to detect subtle deviations in neural dynamics associated with neuropsychiatric conditions (Damaraju et al., 2014; Preti et al., 2017), thus informing targeted interventions and enabling clinicians to monitor treatment efficacy over time. Furthermore, the capacity for real-time adaptation paves the way for advanced brain–computer interfaces that adjust to a user’s cognitive state, optimizing interaction and control in assistive technologies (Kucyi et al., 2024b). Collectively, these advancements underscore the transformative potential of personalized AI models to bridge the gap between artificial and biological intelligence, offering adaptive, interpretable, and human-aligned systems across a wide spectrum of fields.

### v2

@@ Abstract
The integration of human and artificial intelligence offers a powerful avenue for advancing our understanding of information processing, as each system provides unique computational insights. However, despite the promise of human-AI integration, current AI models are largely trained on massive datasets, optimized for population-level performance, lacking mechanisms to align their computations with individual users’ perceptual semantics and neural dynamics. Here we show that integrating human behavioral insights and millisecond scale neural data within a fine tuned CLIP based model not only captures generalized and individualized aspects of perception but also over doubles behavioral performance compared to the unmodified CLIP baseline. By embedding human inductive biases and mirroring dynamic neural processes during training, personalized neural fine tuning improves predictions of human similarity judgments and tracks the temporal evolution of individual neural responses. Our work establishes a novel, interpretable framework for designing adaptive AI systems, with broad implications for neuroscience, personalized medicine, and human–computer interaction.
@@ 1 Introduction
Modern Artificial Intelligence (AI) systems, much like the human brain, emerge from a complex interplay of model architecture, learning rules, objective functions, and training datasets [1, 2]. Yet, despite advances in these components, there remains a persistent gap between AI’s computational processes and the nuanced flexibility inherent in human cognition [3, 4, 5]. Humans effortlessly operate across multiple levels of semantic abstraction, fluidly transitioning between broad categorizations and nuanced distinctions, a capability that even advanced AI models struggle to replicate [7, 8, 9]. Closing this gap by aligning AI more closely with human cognition not only enhances model performance and interpretability but also transforms these systems into powerful tools for probing diverse perceptual processes. To fully harness these benefits, however, AI models must be aligned at the individual level. By capturing the unique cognitive signatures of individual users, personalized AI models offer unprecedented insights into the subtle differences in human perception and provide a novel avenue for understanding the inner workings of individual minds.
@@
While shared architectural principles between deep neural networks (DNNs) and human cortical processing [11] facilitate some degree of brain alignment, recent studies highlight that human-AI alignment may depend more critically on training methodologies than on model architecture itself [12]. Specifically, the choice of training components, such as the objective function significantly influences the learning of human-like invariances [13]. One promising path is therefore to explicitly incorporate human inductive biases into AI systems, guiding their learning toward cognitive patterns found in human perception. Prior research has leveraged Bayesian model selection and contrastive learning to demonstrate that infusing these biases can enhance both human alignment and general machine learning performance [14, 15, 16].
To directly bridge this gap and build on insights that training methodologies critically shape human-AI alignment [12, 13], we fine-tuned the Contrastive Language–Image Pre-training (CLIP) model [17] using human-derived behavioral embeddings generated from large-scale perceptual decision-making datasets [18, 19]. These behavioral embeddings encapsulate interpretable, generalized dimensions of human mental representations, enabling the model to predict human behavioral judgments and neural activity more accurately, significantly surpassing conventional alignment benchmarks [20].
However, while behavioral embeddings capture generalized perceptual dimensions, neural data provide a richer and more granular window into human cognition, revealing millisecond-scale temporal dynamics [21]. Neural data also uniquely allow researchers to more deeply investigate individual variability, especially valuable in populations where behavioral measurements are impractical or challenging, such as infants or non-verbal individuals [49]. To leverage these benefits, we developed a dynamic neural fine-tuning method (CLIP-HBA-MEG), aligning our models with rapidly evolving neural patterns measured via magnetoencephalography (MEG).
By first confirming the robustness of our model at the group-level across diverse stimuli and conditions [19], we extended our approach to fine-tune AI models at the individual participant level. This personalization step allowed us to precisely capture unique neural signatures and cognitive styles, bridging artificial and biological cognition in unprecedented ways. As these personalized models embed distinctive individual perceptual patterns, they not only improve AI’s interpretability and performance but also offer transformative tools for investigating individual differences in cognition. Ultimately, this personalized framework carries deep implications across medicine, cognitive neuroscience, adaptive human-computer interfaces, and tailored AI development, enriching our understanding of both artificial intelligence and the diverse landscape of the human mind.
@@ 2 Results
Current AI models trained on large-scale image datasets develop inductive biases based on the statistical patterns within those datasets. However, a significant gap exists between machine-derived inductive biases and the way humans develop theirs through perceptual experience [23]. To address this, we introduced human-like inductive biases to the CLIP Model in the form of 66 Sparse Positive Similarity Embedding (SPoSE) dimensions, derived from large-scale behavioral studies [24]. These dimensions have been shown to effectively predict human similarity judgments for object distinctions [25].
We developed CLIP-HBA-Behavior (Figure 1) by fine-tuning the CLIP model, using the 66 SPoSE dimensions as textual inputs while simultaneously feeding THINGS images [18] as visual inputs. The CLIP architecture encodes both modalities—textual and visual—separately, before binding them via dot product similarity. This mechanism evaluates the encoded visual features against each of the 66 dimensions, resulting in a 66-dimensional embedding representation for each image input.
@@ 2.1 Fine-Tuning CLIP with Human Behavioral Embeddings Enhances Alignment
As shown in Figure 2A, we evaluated the behaviorally fine-tuned model’s performance on a set of 48 object images from the THINGS database, which were fully sampled in a behavioral odd-one-out task [25]. These 48 objects were excluded from training to prevent data leakage. To assess the alignment between the model’s predictions and human behavior, we calculated the Spearman rank correlation ($\rho$) between the model’s predicted Representational Dissimilarity Matrix (RDM) [26] and the ground-truth behavioral RDM. The behaviorally fine-tuned CLIP-HBA-Behavior achieved a Spearman correlation of 0.78 ($p<10^{-229}$) and a 95% confidence interval (CI) of [0.75, 0.80].
@@ 2.2 Improved Performance on Alignment Benchmark with CLIP-HBA-Behavior
To test our model’s robustness and performance, we benchmarked it against a state-of-the-art human visual similarity judgment metric—NIGHTS (Novel Image Generations with Human-Tested Similarity), a benchmark of 20,019 image triplets with human-tested perceptual similarity scores [20].
@@ 2.3 Improved Neural Alignment with CLIP-HBA-Behavior
While the enhanced behavioral correlation of our model was anticipated due to its training on behavioral data, understanding whether the model’s embedding space truly becomes more human-like requires evaluating how training exclusively on behavioral data impacts its alignment with neural representations. Therefore, we further evaluated the alignment between the model’s representations against neural data obtained from MEG recordings [19].
@@
  • Colored Images without Background (Figure 2B, top right): Clear object stimuli with backgrounds featuring humans, animals, fruits, and man-made objects [27].
  • Monochromatic Clear Images (Figure 2B, bottom left): Clear, monochromatic images of animate and inanimate objects [28].
  • Monochromatic Images with Blur (Figure 2B, bottom right): Blurry, monochromatic images of animate and inanimate objects [28].
@@ 2.4 Dynamic Neural Fine-Tuning Captures Millisecond-Level Neural Dynamics
Building on the success of perceptual fine-tuning with behavioral data, which are large-scale, generalizable, and collected across many trials and participants, we sought to explore whether the model can improve by learning directly from human neural representations. Although typically collected from smaller participant samples, neural data can often be better suited to capture diverse neural profiles across different populations [29]. While behavioral embeddings provide static and generalizable perceptual representations, neural data offer dynamic insights, reflecting millisecond-level temporal patterns of human brain activity that cannot be captured through behavioral measures alone. Furthermore, neural data have the potential to better account for individual variability, providing a more personalized lens into how visual stimuli are processed [30]. To explore these possibilities, we developed a dynamic neural fine-tuning process using MEG signals, as demonstrated in Figure 3. The dynamically fine-tuned model, referred to as CLIP-HBA-MEG, adapts its visual stimulus embeddings to align with distinct stages of visual processing and the emergence of semantic representations.
@@ 2.6 Viewable Dynamic Attention of CLIP-HBA-MEG
Leveraging the dynamic embedding spaces of the neurally aligned CLIP-HBA-MEG, we can visualize the specific regions and pixels the model prioritizes to produce human-like neural responses at a millisecond resolution. By combining randomized input sampling of the visual stimuli [31] [32] with the model’s dynamically learned visual scaler, we can quantify both where and to what extent the model attends within any image following stimulus onset. As shown in Figure 5, the highlighted regions indicate where the model predicts human observers would direct their attention at a given time point.
@@ 2.7 Individualized Models Capture Participant-Specific Neural Dynamics
As shown in Figure 6, we extended the dynamic neural fine-tuning pipeline of CLIP-HBA-Dynamic to train 15 personalized models, each fine-tuned on the individual MEG data of a single participant [11]. These personalized models achieved enhanced neural alignment specific to their respective participants, exhibiting sustained higher temporal alignment compared to the lower-bound noise ceiling, starting around 200 ms after stimulus onset.
@@ 3.1 Scalable and Transferable Applications
Building on these encouraging findings, our first area of focus was the scalability and transferability of the CLIP-HBA framework. CLIP-HBA-MEG leverages dynamic neural signals to capture real-time perceptual processes and higher-order conceptual organization, offering a nuanced window into human perception that static models cannot provide. In contrast to models that focus primarily on behavioral proxies—such as those that linearly map behavioral embeddings into model output [33] or perform generalizable local and global transforms to predict human similarity ratings [34]—CLIP-HBA-MEG establishes a direct relationship with neural activity by explicitly capturing the emergence of information from dynamic neural signals. Recent research has shown that even when behavioral outcomes converge, the underlying neural dynamics can differ substantially between individuals [35, 36]. Such findings underscore the critical importance of capturing dynamic neural activity. While earlier models—such as ReAlnet [4]—attempt to capture cross-modal EEG-fMRI alignments, they lack the temporal resolution necessary to track moment-to-moment perceptual shifts. Moreover, the CLIP-HBA framework can be fine-tuned sequentially, influencing the entire stream of information processing, which uniquely positions it as both a robust model of human perception and a versatile platform that can integrate diverse types of human behavior (e.g., MEG, MRI, behavioral data, electrophysiology). This adaptability is particularly useful in neuroscience applications, where neural data is often limited in scale.
@@ 3.2 Explainable Human-Driven AI Systems
In addition to capturing real-time perceptual processes, CLIP-HBA provides fully interpretable dimensions within the model embedding space [25] that are absent in other models designed to incorporate human inductive biases [4]. By integrating this embedding space with dynamic saliency mapping, we offer millisecond-level insights into both the model’s visual focus and the organization of its internal representations. Furthermore, findings indicate that when model decisions are transparent and systematically quantifiable, it becomes easier to evaluate the consistency of predictions, identify potential biases, and isolate errors for correction [37, 38]. These results support the view that interpretability is essential in applications where accurate, accountable decision-making is critical. Together, these advancements move AI closer to human-like cognition by promoting systems whose internal operations are more accessible and amenable to rigorous oversight [39, 14].
@@ 3.3 Expandable Architectures and Modalities for Future Directions
While our study demonstrates the effectiveness of fine-tuning CLIP to align AI models with human representations using behavioral and MEG data, our approach is not inherently limited to a specific model architecture, type of neural data, or sensory modality. The key factors contributing to the success of our method are the carefully designed loss functions and the modified mechanisms, such as feature reweighting [40] and noise injections. These techniques are broadly applicable to other transformer-based and deep neural network architectures, ensuring that our framework remains adaptable beyond CLIP. This allows our approach to be well-suited for expanding to additional sensory modalities. While our current work focuses on vision, future research could integrate other modalities such as auditory representations for multisensory learning [41]. By fine-tuning models with neural and behavioral data across multiple sensory domains, we can take the first steps toward a truly multisensory AI system—one that more closely mimics the way the human brain integrates information from diverse sensory inputs [42, 43]. This expansion would not only enhance AI’s ability to model human perception but also open new possibilities for multisensory cognitive research and human-AI interaction.
@@ 3.4 Applications of Personalized Models
With the framework’s adaptability established, we now turn to its potential for personalization—a critical aspect of aligning AI with individual human cognition. The CLIP-HBA framework represents a significant advancement in personalized AI by integrating human behavioral embeddings with neural data to tailor models to individual cognitive profiles. This approach is deeply rooted in recent cognitive neuroscience research, which suggests that aligning artificial neural networks with biological cognition can enhance both interpretability and robustness [44]. By fine-tuning on millisecond-level MEG signals, the CLIP-HBA-MEG variant directly optimizes neural alignment, reflecting the neuroconnectionist framework’s emphasis on incorporating biological constraints rather than solely focusing on machine-centric performance. Moreover, the framework’s dynamic adaptation of learned representations based on neural feedback mirrors the functionality of digital twin systems [45]—widely used in industrial and medical applications to model dynamic systems with real-time sensor data—thus acting as a cognitive digital twin that adjusts to an individual’s unique perceptual and neural idiosyncrasies.
Beyond the realm of neuroscience, personalized AI models like CLIP-HBA are making transformative strides in various domains, including education, healthcare, and industry. Advances in AI-driven learning style detection have demonstrated that machine learning can accurately classify users into distinct cognitive subtypes—such as visual, auditory, or kinesthetic learners [46]—which aligns with CLIP-HBA’s ability to fine-tune representations based on individual neural responses. Such personalization is critical for applications where individual differences matter; for example, in clinical contexts, these models can be harnessed to detect subtle deviations in neural dynamics associated with neuropsychiatric conditions [47, 48], thus informing targeted interventions and enabling clinicians to monitor treatment efficacy over time. Furthermore, the capacity for real-time adaptation paves the way for advanced brain–computer interfaces that adjust to a user’s cognitive state, optimizing interaction and control in assistive technologies [49]. Collectively, these advancements underscore the transformative potential of personalized AI models to bridge the gap between artificial and biological intelligence, offering adaptive, interpretable, and human-aligned systems across a wide spectrum of fields.